OpenStudy (tanner23456):

Find the length of the curve y= (x^5)/6 + 1/(10x^3)

OpenStudy (amistre64):

isnt it the integration of sqrt(1+y'^2) ?

OpenStudy (tanner23456):

Yes it is So for the derivative I got \[\frac{ dy }{ dx }=\frac{ 5x^4 }{ 6 } + \frac{ -3x^{-4} }{ 10}\]

OpenStudy (amistre64):

i wonder if doing a polar would be simple or not

OpenStudy (tanner23456):

So, the integral would be I forgot the limits are 1 to 2 \[\int\limits_{1}^{2} \sqrt{1+(\frac{ 5x^4 }{ 6} + \frac{ -3 }{ 10x^4 })^2}dx \] ?

OpenStudy (amistre64):

so far that is to format can you simplify it any?

OpenStudy (tanner23456):

Yeah, that's what I was trying to do. If I did it right I got the derivative to simplify to \[\frac{ 25x^8-9 }{ 30x^4 }\]

OpenStudy (amistre64):

+9 but yes, and that is for x>0

OpenStudy (tanner23456):

I believe it's -9 because the 3 is negative in the derivative.

OpenStudy (tanner23456):

So am I supposed to integrate this now \[\int\limits_{0}^{1}(1+\frac{ (25x^8-9)^2 }{ 900x^8 })^{1/2}dx\]

OpenStudy (tanner23456):

never mind I get the simplification

OpenStudy (amistre64):

then yes, all thats left is to:\[\int\frac{25x^8}{30x^4}+\frac{9}{30x^4}\] then yes, all thats left is to:\[\int\frac{5}{6}x^4+\frac{3}{10}x^{-4}\]